OpenStudy (anonymous):

Which part of the quadratic formula tells you whether the quadratic equation can be solved by factoring? −b b2 − 4ac 2a Use the part of the quadratic formula that you chose above and find its value given the following quadratic equation: 2x2 + 7x + 3 = 0 I can't understand what i need to put as a answer.

OpenStudy (anonymous):

is the answer 1 number or is it a pair?

OpenStudy (anonymous):

as i see it the answer is b2 − 4ac and its value from the equation is 7^2-4*2*3 =25

OpenStudy (anonymous):

what do you mean the value of the equalization?

OpenStudy (anonymous):

equation

OpenStudy (anonymous):

that part of the question Use the part of the quadratic formula that you chose above and find its value given the following quadratic equation: 2x2 + 7x + 3 = 0 the quadaratic function is written as f(x)= ax^2+bx+c so i choosed b2-4ac and from your equation a=2,b=7 and c=3 so b2-4ac is equal to 25

OpenStudy (anonymous):

ok thanks

OpenStudy (anonymous):

if you want to know more about that term (b2-4ac) it is called the Discriminant you can see its page on Wikipedia

OpenStudy (anonymous):

ok